# Energy-Conscious LLM Decoding: Impact of Text Generation Strategies on GPU Energy Consumption

### v1

@@ Abstract
Decoding strategies significantly influence the quality and diversity of the generated texts in large language models (LLMs), yet their impact on computational resource consumption, particularly GPU energy usage, is insufficiently studied. This paper investigates the relationship between text generation decoding methods and energy efficiency, focusing on the trade-off between generation quality and GPU energy consumption across diverse tasks and decoding configurations. By benchmarking multiple strategies across different text generation tasks, such as Translation, Code Summarization, and Math Problem Solving, we reveal how selecting appropriate decoding techniques with their tuned hyperparameters affects text quality and has measurable implications for resource utilization, emphasizing the need for balanced optimization. To the best of our knowledge, this study is among the first to explore decoding strategies in LLMs through the lens of energy consumption, offering actionable insights for designing resource-aware applications that maintain high-quality text generation.
@@ 1 Introduction
In the past few years, we have witnessed a significant increase in the applications of Large Language Models (LLMs). Powerful text generation LLMs such as OpenAI’s ChatGPT [1], Google’s Gemini [2], and Anthropic’s Claude Sonnet [3] have been completely intertwined in our daily lives. The ability of these models to generate high-quality coherent response text is significantly influenced by the choice of decoding strategies during inference. Decoding strategies transform the probability distributions generated by LLMs into fluent and relevant output text [4, 5, 6]. They determine how tokens are selected at each step of text generation. This transformation termed as decoding is as crucial as the model itself in achieving the best quality of the generated text. However, the selection of the appropriate decoding strategy and tuning their specific parameters remain insufficiently explored [4]. In most cases, the evaluations of the newly proposed, state-of-the-art LLMs are only conducted using the default decoding method, instead of studying the effects of adopting different strategies and varying their parameters to achieve the best benchmarking results.
@@
On the other hand, the popularity and increasing applications of LLMs across various domains has caused the AI community to continuously scale the model sizes up to hundreds of billion parameters [7]. This growth has led to a significant surge in computational demands causing a considerable energy consumption associated with the hardware these models are being deployed, especially the GPUs [8, 9]. GPUs have become indispensable to modern high-performance computing (HPC) infrastructure due to their efficient parallel processing [10]. This enhanced capability has made them an excellent accelerator for training and deploying LLMs. However, their significant energy consumption has raised serious alarms recently due to their increasing scale and the global push towards sustainable AI [8, 11]. Although several research have studied the energy consumption and environmental footprints regarding training LLMs [12, 13, 14, 15], the energy consumption concerns associated with LLM inference have been studied much less. This contrasts with the fact that LLM inference has become the main contributor of AI energy footprints recently [16].
@@
Prior research has primarily focused on the introduction of new text generation decoding strategies in the past few years [17, 18, 19, 20, 21]. Each one of these techniques claims to be a superior method based on text quality metrics on specific text generation tasks. While the effects of various decoding strategies on text quality, diversity, and coherence have been studied recently [4, 5], their implications for computational efficiency and energy consumption remain unexplored. Decoding strategies differ in their computational complexity and, consequently, their energy demands. Decoding strategies differ in their computational complexity and, consequently, their energy demands. For instance, beam search, which explores multiple hypotheses simultaneously, is more computationally demanding compared to simpler methods like greedy search. Understanding how these decoding techniques affect energy consumption is necessary for optimizing the LLM inference in resource-constrained devices and achieving an appropriate balance between model performance and sustainability goals.
The main contribution of this research is to analyze the impact of decoding strategies on GPU energy consumption, including:
@@
By providing these contribution, our research aims to contribute to the current literature of compute performance and energy utilization characteristics of LLM inference. Additionally, we hope it encourages further analysis, benchmarking, and broader sharing of systematic performance insights across a wider range of large models, particularly under varying hardware configurations. This can be especially relevant for researchers and users with limited computational resources.
@@ 2.1 Decoding Strategies
2.2 Deterministic Decoding Strategies
@@
  • Beam Search: Instead of greedily selecting the tokens, [23] proposed keeping track of multiple candidate sequences (a beam), expanding them at each generation step, and selecting the top ones based on their overall scores after multiple steps. The number of these candidate sequences is controlled by the beam width, which serves as a hyperparameter.
  • Diverse Beam Search (DBS): [24] introduced an extension of the original Beam Search method by dividing the candidate sequences into multiple subgroups while enforcing diversity within each group. The key hyperparameters for this approach are the beam width and the number of groups.
  • Contrastive Search (CS): [19] introduced a contrastive framework to penalize repetitive token selections while maintaining coherence. After selecting the top-k candidates from the model, they are assessed based on the likelihood assigned by the LLM and a degeneration penalty that evaluates how close the token is relative to the previous context. This method have two hyperparameters: k, which controls candidate selection based on top-k, and $\alpha$, which controls degeneration.
  • DoLa: [20] designed a decoding technique to improve factual accuracy in LLMs. Unlike other decoding methods that rely solely on the final layer’s logits, DoLa compares the output logits of earlier layers with those of the final layer. The specific layers to be contrasted with the final one can be set as a hyperparameter.
  • Assisted Decoding (AD): A novel strategy, also known as speculative decoding, was introduced by [25] to accelerate LLM inference. This technique uses an assistant model to generate multiple candidate tokens in advance, which are then verified by a larger target model. In this way, the number of sequential computations during inference is significantly reduced, speeding up the decoding process. Since we used only the n-gram-based speculative decoding in the greedy setting and not the model-based approach, we categorize it as deterministic. The hyperparameter in this case is the number of tokens to be output as candidate tokens.
2.3 Stochastic Decoding Strategies
@@
  • Top-k: [27] introduced a method which sampling is performed from the $k$ most probable candidates in the vocabulary at each token generation step. A smaller $k$ makes the output more deterministic, whereas a larger k allows tokens with lower probabilities to be considered in the selection process.
  • Top-p (Nucleus): Rather than selecting a fixed number of tokens and sampling from them, [28] proposed selecting tokens based on their probability distribution. They suggested choosing the minimal set of tokens whose cumulative probability goes beyond a probability $p$
  • Epsilon Sampling: [29] introduced another truncation sampling approach by selecting only the tokens whose conditional likelihoods are above an entropy-dependent threshold $\epsilon$. This ensures that the highly unlikely tokens are excluded.
  • Typical Sampling: To generate text that resembles most to human writing, [30] proposed selecting tokens based on information entropy. In the sampling approach, we select tokens based on how close a token’s probability is to the expected information content at each step. Thus, tokens that are too predictable or too unlikely are discarded, with the selection process controlled by a typicality threshold.
  • Min-p Sampling: [21] proposed a dynamic token selection strategy to overcome the limitations of other truncation-based sampling methods. Their approach adjusts the filtering threshold at each step based on the probability of the most likely token. The key hyperparameter is a base probability, which is multiplied by the likelihood of the most probable token to dynamically adjust the threshold for the next generation step.
@@ 3.1 Compute Resources
All the experiments were conducted on a compute node equipped with two NVIDIA A100-PCIE-40GB GPUs and a Dual 64-c ore AMD EPYC 7763 Processor @ 2.45GHz. The system is managed by SLURM [31], an open-source workload manager that facilitates resource allocation, job scheduling, and queue management. In each job (experiment), we requested only one GPU. Before running the main job, we start monitoring and collecting the GPU resource utilization data over the job’s run time. This monitoring is stopped immediately after the main job is finished. The maximum power draw for both GPUs is set to 250 watts.
All the experiments were run on an isolated compute node to minimize the variations in the LLM inference measurements. When running the same experiment (a specific decoding method with fixed parameters with same input data) multiple times alongside other users’ jobs in a compute node, we sometimes observed significant variations in the total inference time and consumed energy. These fluctuations in LLM inference workloads are due to the shared nature of the system (the potential disruptions of other jobs) and distinct compute characteristics of LLMs [16]. Consequently, we submitted all our jobs on isolated node to mitigate the significant variations.
@@ 3.2 Monitoring Tools
We use the nvidia-smi [32] utility to monitor and manage GPUs in our experiments. This command-line utility provides detailed information regarding GPU performance metrics including GPU utilization, temperature, memory usage, memory clock frequency, power consumption, and etc. We use this real-time monitoring utility to record GPU resource usage including utilization and power consumption at 1000 ms intervals for each experiment.
To quantify the total energy consumption, we rely on the $E=\sum_{i=1}^{n}P_{i}\cdot\Delta t_{i}$ formula that relates energy to power ($P_{i}$) and time period ($\Delta t_{i}$). Since we recorded power measurements at 1-second intervals ($\Delta t_{i}=1$), assuming constant power over each second, the energy consumption for each second simplifies to the recorded power (in watts). As a result, the total energy consumption in joules is equivalent to the summation of the recorded power values over the entire inference duration. GPU energy consumption, especially in LLM workloads, results in significantly large values when measured in joules, as joules are a relatively small unit for quantifying energy. To facilitate easier comparison between decoding methods, we converted the energy consumption from joules (J) to watt-hours (Wh) for reporting in this paper($1\mathrm{Wh}=3.6\times 10^{3}\,\mathrm{J}$).
@@ 3.3 Benchmarks
For reproducibility, we used LM-evaluation-harness [33] to benchmark and evaluate the chosen LLMs across different decoding methods. This is an open-source framework designed for evaluating and testing the performance of the LLMs across a wide range of tasks and benchmarks. We evaluated our selected decoding strategies across three text generation tasks: Translation, Code Summarization, and Math. For all three tasks, we utilized only a subset of the full datasets. Using the complete dataset would unnecessarily increase inference time and energy consumption during experimentation, without bringing any meaningful insights to the study. We emphasize that we reported only the available performance metrics for the generated texts as provided by the LM-evaluation-harness framework across our selected tasks.
@@ 3.3.1 Translation
We selected WMT16 [34] dataset to evaluate the translation performance. WMT is a widely recognized competition for machine translation tasks, providing high-quality datasets across multiple languages. In our experiments, we only focused on German-English and English-German and only considered the first 500 input prompts of the whole dataset in both cases. Also, the number of few-shots was set to 0 for all translation jobs.
We report the BLEU [35] in our translation experiments. This widely-used evaluation metric measures the n-gram overlap between the generated output and the ground truth, with higher scores indicating the better generative performance.
3.3.2 Code Summarization
The CodeXGLUE [36] benchmark suit is a collection of datasets designed to evaluate the AI models on various code-related tasks. Specifically, we used Code Summarization [37] task to generate concise and coherent descriptions of source code snippets. This task evaluates the LLM’s ability to understand the semantics of code and translate it into meaningful summaries. We only summarized the first 100 code snippets of Python and Java Script programming languages in to natural language descriptions while the number of few-shots was set to 0 for all experiments.
We use the Smoothed BLEU [38] evaluation metric, a modified version of the standard BLEU, to alleviate the issue of non-overlapping high-order n-grams in the code summarization task [36].
@@ 3.3.3 Math
We utilized the GSM8K dataset [39] to evaluate the reasoning and mathematical problem-solving capabilities of LLMs under different decoding strategies. Although the popular GSM8K consists of 8K grade-school-level math problems, we only focused on the first 100 problems out of the whole benchmark. In this specific task, we used 5-shot prompts in our experiments.
@@
For this benchmark, the Exact Match accuracy is reported as the performance of the generated text compared to the ground truth. As explained in [40], the exact match checks if each token in the a language model’s output aligns with the greedy decoding generation of the input at each step. If all tokens align perfectly, the LLM’s generated text is considered an exact match. Thus, a higher value indicates better text generation performance.
3.4 LLM
In our experiments, we used Qwen 2.5 [41], a recent series of decoder-only LLMs. Specifically, we experimented different decoding strategies with 7B-instruction-tuned model. We selected this model size due to its robust performance across a wide range of text generation tasks and to accelerate the decoding process in our experiments.
3.5 Strategies & Hyperparameters

### v2

@@ Abstract
Decoding strategies significantly influence the quality and diversity of the generated text in Large Language Models (LLMs), yet their impact on computational resources, particularly GPU energy consumption, is insufficiently studied. This paper investigates the relationship between text generation decoding techniques and energy efficiency, focusing on the trade-off between generation quality and GPU energy usage across diverse tasks and decoding configurations. By benchmarking multiple strategies across various tasks, including Translation, Math Problem Solving, Coding, and Open-ended text generation, we reveal how selecting appropriate decoding techniques with their tuned hyperparameters affects text quality and has measurable implications for energy consumption. Our findings show that the choice of decoding strategy can greatly impact GPU energy usage, even when it has a minimal effect on output quality. Different strategies also involve trade-offs between quality and energy efficiency, and no single decoding method is best in all cases across every metric. To the best of our knowledge, this is one of the first studies to examine decoding strategies in LLMs from the perspective of energy consumption, providing useful insights for building energy-efficient applications without compromising text generation quality.
@@ 1 Introduction
In the past few years, we have witnessed a significant increase in the applications of Large Language Models (LLMs). Powerful text generation LLMs such as OpenAI’s ChatGPT [1], Google’s Gemini [2], and Anthropic’s Claude Sonnet [3] have been completely intertwined in our daily lives. The popularity and increasing applications of LLMs across various domains has caused the AI community to continuously scale the model sizes up to hundreds of billion parameters [4]. This growth has led to a significant surge in computational demands causing a considerable energy consumption associated with the hardware these models are being deployed, especially the GPUs [5, 6]. GPUs have become indispensable to modern high-performance computing (HPC) infrastructure due to their efficient parallel processing [7]. This enhanced capability has made them an excellent accelerator for training and deploying LLMs. However, their significant energy consumption has raised serious alarms recently due to their increasing scale and the global push towards sustainable AI [5, 8]. Although most of the AI environmental research have revolved around the energy consumption and carbon footprints during training LLMs [9, 10, 11, 12, 13], the energy concerns associated with LLM inference have been studied much less. Yet, the LLM inference demand is substantially larger [14, 15]. Among recent studies focusing on the energy footprint of LLM inference is an analysis by [16], which measured the energy consumption of GPT-4o single query. Their findings show that earlier research presented by [17] estimating ChatGPT’s energy usage at about 3 watt-hours per query was significantly exaggerated, with the more accurate estimation being closer to 0.3 watt-hours per query. Similarly, a recent study by [18] reported an energy consumption of 0.24 watt-hours per prompt in Google’s Gemini. Similarly, [5] reported the energy usage of different LLaMA model sizes. For the LLaMA-65B model, they reported that with a maximum generation length of 512 tokens, the energy consumption per query is approximately 0.3 watt-hours.
@@
On the other hand, the ability of LLMs to generate high-quality coherent response text is significantly influenced by the choice of decoding strategies during inference. Decoding strategies transform the probability distributions generated by LLMs into fluent and relevant output text [19, 20, 21]. They determine how tokens are selected at each step of text generation. This transformation termed as decoding is as crucial as the model itself in achieving the best quality of the generated text. However, the selection of the appropriate decoding strategy and tuning their specific parameters remain insufficiently explored [19]. In most cases, the evaluations of the newly proposed, state-of-the-art LLMs are only conducted using the default decoding method, instead of studying the effects of adopting different strategies and varying their parameters to achieve the best benchmarking results. Prior research has primarily focused on the introduction of new text generation decoding strategies in the past few years [22, 23, 24, 25, 26]. Each one of these techniques claims to be a superior method based on text quality metrics on specific text generation tasks. While the effects of various decoding strategies on text quality, diversity, and coherence have been studied recently [19, 20], their implications for computational efficiency and energy consumption remain unexplored. Decoding strategies differ in their computational complexity and, consequently, their energy demands. For instance, beam search, which explores multiple hypotheses simultaneously, is more computationally demanding compared to simpler methods like greedy search. Understanding how these decoding techniques affect energy consumption is necessary for optimizing the LLM inference in resource-constrained devices and achieving an appropriate balance between model performance and sustainability goals. In this respect, the main contribution of this research is to analyze the impact of decoding strategies on GPU energy consumption, including:
@@
  • Distinguishing between decoding latency and GPU energy consumption by analyzing their relationship
Through these contributions, our work adds to the growing literature on the generative performance and energy characteristics of LLM inference. We also hope it motivates further research, benchmarking, and open sharing of systematic performance insights across a broader range of models and hardware settings. Such efforts are particularly valuable given the growing concerns about the energy consumption and environmental impact of large-scale AI models, emphasizing the need for developing more efficient inference methods.
@@ 2 Background
Deterministic Strategies
Deterministic decoding prioritizes maximizing the probability of output sequences according to the model’s learned distribution [24]. Although these strategies lead to higher generative accuracy and coherence, they also limit diversity and creativity, which can result in text degeneration [27, 28]. In this study, we examine a selection of widely adopted deterministic methods:
@@
  • Beam Search: Instead of greedily selecting the tokens, Freitag et al. [30] proposed keeping track of multiple candidate sequences (a beam), expanding them at each generation step, and selecting the top ones based on their overall scores after multiple steps. The number of these candidate sequences is controlled by the beam width, which serves as a hyperparameter.
  • Diverse Beam Search (DBS): Vijayakumar et al. [31] introduced an extension of the original Beam Search method by dividing the candidate sequences into multiple subgroups while enforcing diversity within each group. The key hyperparameters for this approach are the beam width and the number of groups.
  • Contrastive Search (CS): Su et al. [24] introduced a contrastive framework to penalize repetitive token selections while maintaining coherence. After selecting the top-k candidates from the model, they are assessed based on the likelihood assigned by the LLM and a degeneration penalty that evaluates how close the token is relative to the previous context. This method have two hyperparameters: k, which controls candidate selection based on top-k, and $\alpha$, which controls degeneration.
  • DoLa: Chuang et al. [25] designed a decoding technique to improve factual accuracy in LLMs. Unlike other decoding methods that rely solely on the final layer’s logits, DoLa compares the output logits of earlier layers with those of the final layer. The specific layers to be contrasted with the final one can be set as a hyperparameter.
  • Assisted Decoding (AD): A novel strategy, also known as speculative decoding, was introduced by Leviathan et al. [32] to accelerate LLM inference. This technique uses an assistant model to generate multiple candidate tokens in advance, which are then verified by a larger target model. In this way, the number of sequential computations during inference is significantly reduced, speeding up the decoding process. Since we used only the n-gram-based speculative decoding in the greedy setting and not the model-based approach, we categorize it as deterministic. The hyperparameter in this case is the number of tokens to be output as candidate tokens.
Stochastic Strategies
Stochastic decoding introduces randomness in the text generation process to help prevent text degeneration. These methods increase the model’s ability to generate diverse outputs, which is especially useful for creative writing. While these sampling-based methods help increase diversity, the randomness in these approaches can negatively affect the coherence of the generated text [24]. This work considers the following stochastic decoding methods for evaluation:
@@
  • Top-k: Fan et al. [27] introduced a method which sampling is performed from the $k$ most probable candidates in the vocabulary at each token generation step. A smaller $k$ makes the output more deterministic, whereas a larger k allows tokens with lower probabilities to be considered in the selection process.
  • Top-p (Nucleus): Rather than selecting a fixed number of tokens and sampling from them, Holtzman et al. [28] proposed selecting tokens based on their probability distribution. They suggested choosing the minimal set of tokens whose cumulative probability goes beyond a probability $p$
  • Epsilon Sampling: Hewitt et al. [34] introduced another truncation sampling approach by selecting only the tokens whose conditional likelihoods are above an entropy-dependent threshold $\epsilon$. This ensures that the highly unlikely tokens are excluded.
  • Typical Sampling: To generate text that resembles most to human writing, Meister et al. [35] proposed selecting tokens based on information entropy. In the sampling approach, we select tokens based on how close a token’s probability is to the expected information content at each step. Thus, tokens that are too predictable or too unlikely are discarded, with the selection process controlled by a typicality threshold.
  • Min-p Sampling: Nguyen et al. [26] proposed a dynamic token selection strategy to overcome the limitations of other truncation-based sampling methods. Their approach adjusts the filtering threshold at each step based on the probability of the most likely token. The key hyperparameter is a base probability, which is multiplied by the likelihood of the most probable token to dynamically adjust the threshold for the next generation step.
@@ 3.1 Compute Resources
All the experiments were conducted on a compute node equipped with two NVIDIA A100-PCIE-40GB GPUs and a Dual 64-core AMD EPYC 7763 Processor @ 2.45GHz. The GPUs have the Thermal Design Power (TDP) of 250 watts. The system is managed by SLURM [36], an open-source workload manager that facilitates resource allocation, job scheduling, and queue management. In each job (experiment), we requested only one GPU. Before running the main job, we start monitoring and collecting the GPU resource utilization data over the job’s run time. This monitoring is stopped immediately after the main job is finished.
All the experiments were run on an isolated compute node to consider the variations in the LLM inference measurements and to minimize the external interference. When running the same experiment (a specific decoding method with fixed parameters with same input data) multiple times alongside other users’ jobs in a compute node, we sometimes observed significant variations in the total inference time and consumed energy. These fluctuations in LLM inference workloads are due to the shared nature of the system (the potential disruptions of other jobs) and distinct compute characteristics of LLMs [37]. Consequently, we submitted all our jobs on isolated node to mitigate the significant variations.
@@ 3.2 Monitoring Tools
We use Nvidia built-in power sensors to measure GPU power consumption in our experiments. Nvidia offers two on-board monitoring tools: Nvidia System Management Interface[38](nvidia-smi) and Nvidia Data Center GPU Manager[39] (dcgm). Nvidia-smi is a command-line utility to monitor and manage Nvidia GPUs. It provides detailed information about GPU performance metrics including utilization, temperature, memory usage, memory clock frequency, power consumption, and etc. While this utility is useful for monitoring, it offers only a limited set of metrics [40]. In contrast, dcgm provides more fine-grained insights than nvidia-smi. It can capture detailed hardware utilization and an easier GPU management in data centers. Both tools allow us to record power consumption at a sampling frequency in units of milliseconds.
3.2.1 Energy Measurement Procedure
We conducted a thorough investigation of what granularity of power measurements works best in our experiments. According to the dcgm documentation, the minimum sampling frequency supported is 100 ms, but the nvidia-smi documentation does not provide any information about its minimum sampling interval. We tested sampling frequencies of 10, 100, and 1000 ms with nvidia-smi. While finer-grained measurements would be ideal, we observed important technical limitations when using on-board monitoring tools. With nvidia-smi, shorter sampling intervals led to significant information loss compared to the 1000 ms interval. For example, at a requested 100 ms interval, we collected only about 9 samples per second instead of the expected 10, indicating a 10% information loss. This issue was even more pronounced at 10 ms, where we gathered only about 70 samples per second instead of 100, representing a 30% loss. These results show that nvidia-smi is less reliable for power monitoring at small sampling intervals. These technical limitations of nvidia-smi, due to its internal mechanisms, have also been reported by Yang et al. [41] and Zou et al. [42].
We also tested sampling intervals of 1000 and 100 ms with the NVIDIA dcgm tool. At the minimum frequency of 100 ms, we observed much less sampling loss compared to the same case with nvidia-smi, showing that dcgm is a better monitoring tool for collecting fine-grained power measurements. When comparing the aggregated energy consumption across different decoding strategies on the watt-hour scale, we found only small differences between nvidia-smi (with a 1000 ms interval) and dcgm (with a 100 ms interval). Importantly, the relative rankings of decoding strategies remained consistent across both tools. However, to better capture momentary power draw fluctuations during LLM inference and instantaneous spikes in GPU utilization, we collected fine-grained power measurements using the dcgm tool with a 100 ms sampling interval.
To quantify the total energy consumption, we rely on the $E=\sum_{i=1}^{n}P_{i}\cdot\Delta t_{i}$ formula that relates energy to power ($P_{i}$) and time period ($\Delta t_{i}$). Since we recorded power measurements at 100-millisecond intervals ($\Delta t_{i}=0.1s$), assuming constant power (in watts) over each sampling period, the total energy consumption in joules is equivalent to the summation of the recorded power values over the entire inference duration multiply 0.1. GPU energy consumption, especially in LLM workloads, results in significantly large values when measured in joules, as joules are a relatively small unit for quantifying energy. To facilitate easier comparison between decoding methods, we converted the energy consumption from joules (J) to watt-hours (Wh) for reporting in this ($1\mathrm{Wh}=3.6\times 10^{3}\,\mathrm{J}$).
@@ 3.3 Benchmarks
For reproducibility, we used LM-evaluation-harness [43] to benchmark and evaluate the chosen LLMs across different decoding methods. This is an open-source framework designed for evaluating and testing the performance of the LLMs across a wide range of tasks and benchmarks. We evaluated our selected decoding strategies across various text generation tasks: Translation, Coding, Math, and Open-ended Generation. For all tasks, we utilized only a subset of the full datasets. Using the complete dataset would unnecessarily increase inference time and energy consumption during experimentation, without bringing any meaningful insights to the study. We emphasize that we reported only the available performance metrics for the generated texts as provided by the LM-evaluation-harness framework (v0.4.8) across our selected tasks.
@@ 3.3.1 Translation
We selected the WMT16 [44] dataset to evaluate the translation performance. WMT is a widely recognized competition for machine translation tasks, providing high-quality datasets across multiple languages. In our experiments, we focused on German-English and English-German and only considered the first 500 input prompts of the whole dataset in both cases. For evaluation, We reported the BLEU [45] metric. This widely-used evaluation metric measures the n-gram overlap between the generated output and the ground truth, with higher scores indicating the better generative performance.
@@ 3.3.2 Math
We utilized the GSM8K [46] dataset to evaluate the reasoning and mathematical problem-solving capabilities of LLMs under different decoding strategies. Although this popular dataset consists of 8K grade-school-level math problems, we only focused on the first 100 problems out of the whole benchmark. In this specific task, we used 5-shot prompts in our experiments. For this benchmark, the Exact Match accuracy is reported as the performance of the generated text compared to the ground truth. This metric checks if each token in the a language model’s output aligns with the greedy decoding generation of the input at each step [47]. If all tokens align perfectly, the LLM’s generated text is considered an exact match. Thus, a higher value indicates better text generation performance.
3.3.3 Coding
The HumanEval [48] benchmark suit is a collection of programming problems designed to evaluate the AI models on code generation capabilities. The benchmark asses an LLM’s ability in code synthesis tasks, typically measured using the pass@k metric. We evaluated our models on 100 programming problems in the benchmark and reported the pass@1 accuracy.
3.3.4 Open-ended Generation
We also evaluated our models on the WikiText [49] benchmark, a large dataset of verified articles on Wikipedia curated for open-ended text generation. Wikitext is a standard benchmark for evaluating LLMs in free-form language modeling. We utilized 400 randomly selected articles from the pre-processed dataset provided by Su et al. [50] in our experiments. To evaluate WikiText benchmark, we used MAUVE [51] metric. It quantifies how similar the distribution of generated text is to that of human-written reference text. The resulting score reflects the overlap between the two distributions.
Table 1: Decoding Strategies and Selected Hyperparameters
strategies hyperparameter value
min_p {0.02, 0.05, 0.1, 0.3, 0.5, 0.7, 0.9, 1.0}
epsilon {0.0003, 0.0006, 0.0009, 0.002, 0.004}
contrastive $\alpha:${0.2, 0.4, 0.6} k:k:{5, 25}}
temperature {0.1, 0.3, 0.5, 0.7, 0.9, 1.2}
diverse beam {(4,2), (5,5), (10,2), (10,5)}
top_p {0.3, 0.5, 0.7, 0.8, 0.95}
top_k {5, 25, 50, 100}
typical {0.5, 0.8, 0.9, 0.95}
dola {low, high}
assisted {2, 5, 10}
beam search {2, 5, 10}
LLMs
In our experiments, we used Qwen 2.5 [52] and LLaMA 3.1 [53], two state-of-the-art family of open-source models. Specifically, we experimented with the 7B and 14B instruction-tuned variants of Qwen 2.5, as well as the 8B instruction-tuned variant of LLaMA 3.1. By using models of different sizes and from different model families, our goal is to analyze how architectural choices and parameter scaling influence decoding behavior, runtime, and energy consumption in different inference tasks.
Strategies & Hyperparameters